OpenStudy (anonymous):

Given the points (10, 4) and (7, 0) find the slope. @faisalalif1999

OpenStudy (anonymous):

this is easy

OpenStudy (anonymous):

use slope formula

OpenStudy (vshiroky):

(y2-y1)=(x2-x1)

OpenStudy (anonymous):

\[\frac{ 0-4 }{ 7-10 }\]

OpenStudy (anonymous):

-4,-3 ?

OpenStudy (anonymous):

\[\frac{ -4 }{-3 } = \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

Given the points (-10, 4) and (0, -6) find the slope

OpenStudy (anonymous):

this one I do the same than

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got -10/-10

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

no -10/10 = -1

OpenStudy (anonymous):

-10/-10 = +1

OpenStudy (anonymous):

your slope is -1

OpenStudy (anonymous):

alrigt et me do this one

OpenStudy (anonymous):

Find the slope of the line that goes through the points (4,-7) and (12,9). Slope, m =

OpenStudy (anonymous):

I got 8/16

OpenStudy (anonymous):

@faisalalif1999

OpenStudy (anonymous):

is it good ?

OpenStudy (anonymous):

yes, simplified it is m = 2

OpenStudy (anonymous):

now, i gotta log out. see ya